# Human-LLM Coevolution: Evidence from Academic Writing

### v1

@@ Abstract
With a statistical analysis of arXiv paper abstracts, we report a marked drop in the frequency of several words previously identified as overused by ChatGPT, such as “delve”, starting soon after they were pointed out in early 2024. The frequency of certain other words favored by ChatGPT, such as “significant”, has instead kept increasing. These phenomena suggest that some authors of academic papers have adapted their use of large language models (LLMs), for example, by selecting outputs or applying modifications to the LLM-generated content. Such coevolution and cooperation of humans and LLMs thus introduce additional challenges to the detection of machine-generated text in real-world scenarios. Estimating the impact of LLMs on academic writing by examining word frequency remains feasible, and more attention should be paid to words that were already frequently employed, including those that have decreased in frequency.
@@ 1 Introduction
The detection of machine-generated text (MGT) has also attracted a lot of attention (Tang et al., 2024), but the performance of detectors has also been questioned early on (Sadasivan et al., 2023; Weber-Wulff et al., 2023; Ghosal et al., 2023). Recent studies continue to show that some methods are not sufficiently robust (Zhang et al., 2024b; Wang et al., 2024; Creo and Pudasaini, 2025). The effectiveness of MGT detectors is also related to the model of LLMs and the type of text (Liu et al., 2024), and their accuracy may also be exaggerated (Doughman et al., 2024). The situations likely to arise in reality are more complicated and are not limited to a binary classification framework Zhang et al. (2024a). Thus, examining and analyzing the ongoing evolution of word usage remains a useful and meaningful task.
Figure 1: Frequency in arXiv abstracts of some of the words that were singled out as favored by ChatGPT around April 2024.
@@
Figure 1 illustrates the evolution in the frequency usage of some of the words that were singled out as favored by ChatGPT around April 2024. The frequency of “significant” and “additionally” continues to grow, while that of “is” and “are” continues its declining trend, as noted by Geng and Trotta (2024). Meanwhile, the frequency of some other words (e.g., “intricate” and “delve”) associated with LLMs begins to decrease after March and April 2024, which corresponds to the time when researchers identified these words in AI conference peer reviews (Liang et al., 2024a) and academic papers (Liang et al., 2024b).
Changes in the words used in academic writing, as discussed above, serve as an excellent example of AI and human coevolution. Researchers are constantly proposing new detection techniques, but the language and expressions of LLM users are also likely evolving due to their use of LLMs (Geng et al., 2024). Given the lack of a precise definition, LLM-generated text might even be undetectable in individual instances. Therefore, statistically measuring the impact of LLMs over a large corpus of texts is a more practical option.
@@ 3 Word Frequency Analysis
We compared the results with the abstracts of withdrawn papers, as illustrated in Figures 2(e) and 2(f). Given the small number of withdrawn papers, the 12-month rolling averages of their word frequency are used in the graphs. The frequency of some words, such as “intricate”, is higher in the withdrawn papers, but the difference is not very large, as is also the case in Figures 7(b), 7(c) and 7(d) of the Appendix.
@@ 4 Challenges in Machine-Generated Text Detection
The first 1000 arXiv papers submitted each year from 2018 to 2025 were utilized for this part of the analysis. We also used the following two simple prompts to examine the differences between original arXiv abstracts and those revised by GPT-4o-mini:
@@
The results in Figure 4 reinforce the point that the frequency of certain words increases after LLMs revision. Using prompt P2, aimed at suppressing them, reduces the frequency of such words, although it does not eliminate them completely.
Figure 4: Comparison of word frequencies before and after LLM processing (with prompt P1 or P2).
@@
Figure 5: MGT detection results for real and LLM-processing abstracts (with prompt P1 or P2). A lower score indicates a greater probability that the text is machine-generated.
@@
5 Conclusions and Discussions
Humans and LLMs are coevolving and we can already conclude that, for this reason, the impact of LLMs on academic writing will fully assert itself over the long term. According to recent studies, people who frequently use ChatGPT for writing tasks can accurately distinguish AI-generated text (Russell et al., 2025), which implies they are also able to foil MGT detectors.
@@ 5 Conclusions and Discussions
Therefore, using the frequency of more common words (Geng and Trotta, 2024) to measure the impact of LLMs on a vast number of publications will be more reliable, although this approach is less suitable for the precise detection of short texts.

### v2

@@ Abstract
With a statistical analysis of arXiv paper abstracts, we report a marked drop in the frequency of several words previously identified as overused by ChatGPT, such as “delve”, starting soon after they were pointed out in early 2024. The frequency of certain other words favored by ChatGPT, such as “significant”, has instead kept increasing. These phenomena suggest that some authors of academic papers have adapted their use of large language models (LLMs), for example, by selecting outputs or applying modifications to the LLM outputs. Such coevolution and cooperation of humans and LLMs thus introduce additional challenges to the detection of machine-generated text in real-world scenarios. Estimating the impact of LLMs on academic writing by examining word frequency remains feasible, and more attention should be paid to words that were already frequently employed, including those that have decreased in frequency due to LLMs’ disfavor.
@@ 1 Introduction
The detection of machine-generated text (MGT) has also attracted a lot of attention (Tang et al., 2024; Chowdhury et al., 2024; Wang et al., 2025), but the performance of detectors has also been questioned early on (Sadasivan et al., 2023; Weber-Wulff et al., 2023; Ghosal et al., 2023). Recent studies continue to show that some methods are not sufficiently robust (Zhang et al., 2024b; Wang et al., 2024; Creo and Pudasaini, 2025). The effectiveness of MGT detectors is also related to the model of LLMs and the type of text (Liu et al., 2024), and their accuracy may also be exaggerated (Doughman et al., 2024). The situations likely to arise in reality are more complicated and are not limited to a binary classification framework Zhang et al. (2024a). Thus, examining and analyzing the ongoing evolution of word usage remains a useful and meaningful task.
Figure 1: The frequency evolution of some words in arXiv abstracts (they were singled out around April 2024 as either favored or disfavored by ChatGPT).
@@
Figure 1 illustrates the evolution in the frequency usage of some of the words that were singled out as favored or disfavored by ChatGPT around April 2024. The frequency of “significant” and “additionally” continues to grow, while that of “is” and “are” continues its declining trend, as noted by Geng and Trotta (2024). Meanwhile, the frequency of some other words (e.g., “intricate” and “delve”) associated with LLMs begins to decrease after March and April 2024, which corresponds to the time when researchers identified these words in AI conference peer reviews (Liang et al., 2024a) and academic papers (Liang et al., 2024b).
Changes in the words used in academic writing, as discussed above, serve as an excellent example of AI and human coevolution. Researchers are constantly proposing new detection techniques, but the language and expressions of LLM users are also likely evolving due to their use of LLMs (Geng et al., 2024). Given the lack of a precise definition, LLM-generated text might be undetectable in certain individual instances. Therefore, statistically measuring the impact of LLMs over a large corpus of texts is a more practical option.
@@ 3 Word Frequency Analysis
We compared the results with the abstracts of the withdrawn papers, as illustrated in Figures 2(e) and 2(f). Given the small number of withdrawn papers, the 12-month rolling averages of their word frequency are used in the graphs. The frequency of some words, such as “intricate”, is higher in the withdrawn papers, but the difference is not very large, as is also the case in Figures 7(b), 7(c) and 7(d) of the Appendix.
@@ 4 Challenges in Machine-Generated Text Detection
[…]
@@
The results in Figure 4 reinforce the point that the frequency of certain words increases after LLMs revision. Using prompt P2, aimed at suppressing them, reduces the frequency of such words, although it does not completely eliminate them.
Figure 4: Comparison of word frequencies before and after LLM processing (with prompts P1 or P2).
@@
Figure 5: MGT detection results for real and LLM-processing abstracts (with prompts P1 or P2). A lower score indicates a greater probability that the text is machine-generated.
@@
5 Conclusion and Discussion
Humans and LLMs are coevolving and we can already conclude that, for this reason, the impact of LLMs on academic writing will fully assert itself over the long term. According to recent studies, people who frequently use ChatGPT for writing tasks can accurately distinguish AI-generated text (Russell et al., 2025), which implies that they are also able to foil MGT detectors.
@@ 5 Conclusion and Discussion
Therefore, using the frequency of more common words to measure the impact of LLMs on a vast number of publications will be more reliable, although this approach is less suitable for the precise detection of short texts.